# ColorVideoVDP: A visual difference predictor for image, video and display distortions

### v1

@@
Abstract
ColorVideoVDP is a video and image quality metric that models spatial and temporal aspects of vision, for both luminance and color. The metric is built on novel psychophysical models of chromatic spatiotemporal contrast sensitivity and cross-channel contrast masking. It accounts for the viewing conditions, geometric, and photometric characteristics of the display. It was trained to predict common video streaming distortions (e.g. video compression, rescaling, and transmission errors), and also 8 new distortion types related to AR/VR displays (e.g. light source and waveguide non-uniformities). To address the latter application, we collected our novel XR-Display-Artifact-Video quality dataset (XR-DAVID), comprised of 336 distorted videos. Extensive testing on XR-DAVID, as well as several datasets from the literature, indicate a significant gain in prediction performance compared to existing metrics. ColorVideoVDP opens the doors to many novel applications which require the joint automated spatiotemporal assessment of luminance and color distortions, including video streaming, display specification and design, visual comparison of results, and perceptually-guided quality optimization.
Figure 1: ColorVideoVDP  predicts the visibility of distortions for a pair of test and reference videos (or images) as seen on a display with a provided specification. The predictions are represented as a single quality value in Just-Objectionable-Difference (JOD) units, a distortion map video, and a distogram, which visualizes the distortions over time, separately for each channel and spatial frequency band.
Keywords: image quality, video quality, visual difference predictor, contrast sensitivity, visual metric
1 Introduction
@@
Accurate reproduction of contrast and color is of central importance to the quality of displayed content. Achromatic artifacts stemming from graphics pipelines, such as visible blur or contrast loss, can be modelled by existing luminance-only metrics such as SSIM Wang et al. [2003] or FovVideoVDP Mantiuk et al. [2021], but color ones, such as chroma subsampling, cannot. On the other hand, color difference formulas, such as the popular CIEDE2000 Sharma et al. [2005]; CIE [2018], do not model spatial or temporal aspects of vision, and as a consequence may ignore important aspects of an artifact, such as the spatial distribution or changes of color distortion over time. Notably, spatiotemporal color artifacts are especially problematic in modern display applications, in particular for emerging display technologies such as wide-color-gamut, virtual and augmented reality (XR) displays. The latter require novel architectural solutions, which lead to chromatic artifacts like color fringing caused by lens aberrations, or chromatic nonuniformity due to optical waveguides (see Fig. 1).
@@
This work presents ColorVideoVDP  [^1] The code for the metric is included in the supplementary materials and will be released., a full-reference quality metric that models spatiotemporal achromatic and chromatic vision. The metric is built on novel psychophysical models of chromatic spatiotemporal contrast sensitivity and cross-channel contrast masking. Thanks to its psychophysical foundations, the metric accounts for the physical specification of a display (size, resolution, color characteristic) and viewing conditions (viewing distance and ambient light). This is the first video and image quality metric that explicitly models human spatiotemporal and chromatic vision simultaneously and is capable of modelling XR display artifacts.
@@
The key challenge of developing any new quality metric is its effective calibration and robust validation. To that end, and to ensure that the metric can provide reliable predictions for display applications, we collected a new XR-Display-Artifact-Video quality dataset (XR-DAVID) with 8 common display artifacts [^2] The dataset will be released on acceptance, subject to institutional approval. (Sec. 4). To ensure the diversity of distortion and content types, we combined our new XR-DAVID dataset with a large HDR/SDR image dataset UPIQ Mikhailiuk et al. [2022]. As the combined datasets consist of terabytes of data, calibration required a non-trivial mixture of end-to-end, and feature-space training. This effort allowed us to match and exceed the state-of-the-art results on unseen datasets (in a cross-dataset validation) and on the testing portion of the training datasets (cross-content validation). In Sec. 6, we demonstrate new applications of ColorVideoVDP , including analysis of chroma subsampling, display color tolerance specification, and quantifying observer metamerism variations on a target dataset.
@@
ColorVideoVDP  lacks higher-level models of saliency or annoyance, resulting in lower accuracy when the semantic content has a strong influence on the quality judgments. It was not trained to predict accurate distortion maps Ye et al. [2019] (as no such data is available for video). ColorVideoVDP  does not model the effect of glare (inter-ocular light scatter, found in HDR-VDP Mantiuk et al. [2023]), gaze-contingent vision (found in FovVideoVDP Mantiuk et al. [2021]), eye motion Laird et al. [2006]; Denes et al. [2020], or binocular vision Didyk et al. [2011].
2 Related Work
Table 1: Quality/difference metrics and their capabilities. Columns indicate whether the metric models spatial vision, temporal vision, color vision, and whether it accounts for display geometry and photometry.
Metric Spatial Temporal Color
Display model
Approach
PSNR No No No No
Signal quality
CIEDE2000 CIE [2018] No No Yes Yes
Color difference formula
$\Delta\text{E}_{\text{ITP}}$ ITU-R BT. 2124 [2019] No No Yes Yes
Color difference formula
sCIELAB Zhang and Wandell [1997] Yes No Yes Yes
CSF + Color difference formula
$\Delta\text{E}_{\text{ITP}}^{\text{S}}$ Choudhury et al. [2021] Yes No Yes Yes
CSF + Color difference formula
MS-SSIM Wang et al. [2003] Yes No No No
Multi-scale structural similarity
FSIMc Zhang et al. [2011] Yes No Yes No
Similarity of phase congruency and gradients
VSI Zhang et al. [2014] Yes No Yes No
Saliency + SSIM
LPIPS Zhang et al. [2018] Yes No Yes No
Difference of CNN features
FLIP Andersson et al. [2020] Yes No Yes No
CSF + Color difference + edge detectors
IQT Cheon et al. [2021] Yes No Yes No
CNN features + transformer autoencoder
STRRED Soundararajan and Bovik [2012] Yes Yes No No
Entropy differences in wavelet subbands
VMAF Li et al. [2016a] Yes Yes No No
Features + SVR
HDR-VDP-3 Mantiuk et al. [2023] Yes No No Yes
Psychophysical model
FovVideoVDP Mantiuk et al. [2021] Yes Yes No Yes
Psychophysical model
ColorVideoVDP  (ours) Yes Yes Yes Yes
Psychophysical model
@@ 2 Related Work
[…]
@@
The spatial component of vision was included in a spatial extension of the CIELAB difference formula — sCIELAB Zhang and Wandell [1997]. The authors proposed to compute the CIE $\Delta\text{E}^{*}_{ab}$ color differences on images prefiltered by a contrast sensitivity function (CSF). Flip and HDR-Flip metrics Andersson et al. [2020, 2021] improve on sCIELAB by employing a color difference formula that better quantifies large color differences. Both metrics also emphasize differences at edges, which tend to be more salient. Choudhury et al. proposed to compute color differences in the ITP color space Choudhury et al. [2021], which is more suitable for HDR color values. The strength of such spatial extensions is their simplicity. The main weakness is that such an application of the CSF is overly simplistic — it does not account for the changes in contrast sensitivity with luminance and does not account for supra-threshold vision (e.g. contrast masking and contrast constancy).
@@
The most popular image quality metrics, such as SSIM or MS-SSIM Wang et al. [2003], do not attempt to explicitly model human vision, but, instead, they combine hand-crafted statistical measures that are likely to correlate with quality judgments. Although the early metrics, such as SSIM and MS-SSIM, operate only on the luma channel of the image, some later metrics, including FSIMc Zhang et al. [2011] and VSI Zhang et al. [2014], separate images into luma and two chroma channels, akin to the color space transforms used in video compression. Those metrics, however, do not account for the geometry of a display (e.g. resolution, size, viewing distance), nor for its photometry (e.g. peak brightness, black level). The latter shortcoming can be addressed by employing the Perceptually Uniform (PU) transform Mantiuk and Azimi [2021], which also extends these metrics to operate on high-dynamic-range images.
Zhang et al. Zhang et al. [2018] observed that the activation layers of many convolution neural networks (CNNs) provide features that are well correlated with human judgments of image similarity. Their proposed metric, LPIPS, became very popular in computer vision, despite its underwhelming performance on image quality datasets Ding et al. [2021]. Prashnani et al. [2018] proposed training a CNN on triplets of patches, two distorted and one reference, with supervision based on the Bradley-Terry pairwise comparison model. Akin to LPIPS training, their PieAPP metric was trained on a large dataset of patches with pairwise comparison labels. Cheon et al. [2021] combined CNN features with transformer-based embeddings to regress quality scores. Their IQT metric was ranked first among 13 participants in the NTIRE 2021 perceptual image quality assessment challenge.
@@
Video quality metrics combine both spatial and temporal features. STRRED Soundararajan and Bovik [2012] compares the per-frame entropy of fitted local distributions of wavelet coefficients. This entropy is then weighted by local spatial and temporal variance. VMAF Li et al. [2016a] combines two spatial features (VIF and DLM) with a mean absolute difference of consecutive frames, and then maps those into quality scores using a support vector regression. The temporal processing of both of these state-of-the-art metrics is rather limited as it considers just two consecutive frames and it does not model the temporal characteristics of human vision. Our metric explicitly models achromatic and chromatic visual channels to address this shortcoming.
@@
The visual difference predictors (VDPs), such as DCTune Watson [1993], VDP Daly [1993], HDR-VDP-2 Mantiuk et al. [2011] and HDR-VDP-3 Mantiuk et al. [2023], explicitly model aspects of low-level human vision, such as contrast sensitivity and masking. The advantage of this approach is that these metrics are built on sound psychophysical models and generalize more easily to unseen conditions, such as displays of varying size, resolution, or peak luminance. While it is tempting to assume that this type of general modeling would perform worse or be more computationally expensive than metrics relying on hand-crafted features, as we will demonstrate in Sec. 5, this is not the case. Modern VDPs perform on par or better than metrics with hand-crafted features. When they are optimized to run on a GPU, they are as fast as feature-based metrics.
@@
Our ColorVideoVDP  metric takes inspiration from and is based on similar building blocks as FovVideoVDP Mantiuk et al. [2021] but with several important improvements. First, ColorVideoVDP  models the visibility of chromatic (color) differences by employing a novel spatiotemporal-chromatic contrast sensitivity function (castleCSF) Anonymous [2024]. This involves modeling both spatial and temporal chromatic vision. Second, we account for supra-threshold color differences so that the perceived magnitude of achromatic and chromatic contrast is properly mapped to the metric response. Third, ColorVideoVDP  models within-channel and cross-channel masking, where each of the four modeled channels (two achromatic and two chromatic) can be masked by the combination of contrast in the other channels. Finally, ColorVideoVDP  is trained and tested on multiple image and video datasets, both SDR and HDR, including a novel dataset of display artifacts described in Sec. 4. Unlike FovVideoVDP, we do not model foveated vision due to the increased overhead of modeling interactions between eccentricity and color. While modeling foveated vision is necessary in some specific usage scenarios, such as foveated rendering, it is less useful in most general-use cases, such as video streaming, display engineering, and non-foveated algorithm design.
3 Color video visual difference predictor
@@
Figure 2: Processing stages of ColorVideoVDP. Test and reference images are first processed by the same pipeline: the display model maps pixel values to linear color (CIE 1931 XYZ color space), linear color is transformed to the opponent color space (DKL), the achromatic channel is decomposed into sustained (continuous lines) and transient (dashed lines) temporal channels, then each of those is decomposed into multiple spatial bands (Laplacian pyramid). The decomposed video/image goes into the contrast sensitivity and masking models, explained in more detail in Figure 6. The result of the masking model is pooled across all spatial bands, temporal and color channels, and finally regressed to a JOD score.

### v2

@@
Abstract.
ColorVideoVDP is a video and image quality metric that models spatial and temporal aspects of vision for both luminance and color. The metric is built on novel psychophysical models of chromatic spatiotemporal contrast sensitivity and cross-channel contrast masking. It accounts for the viewing conditions, geometric, and photometric characteristics of the display. It was trained to predict common video-streaming distortions (e.g., video compression, rescaling, and transmission errors) and also 8 new distortion types related to AR/VR displays (e.g., light source and waveguide non-uniformities). To address the latter application, we collected our novel XR-Display-Artifact-Video quality dataset (XR-DAVID), comprised of 336 distorted videos. Extensive testing on XR-DAVID, as well as several datasets from the literature, indicate a significant gain in prediction performance compared to existing metrics. ColorVideoVDP opens the doors to many novel applications that require the joint automated spatiotemporal assessment of luminance and color distortions, including video streaming, display specification, and design, visual comparison of results, and perceptually-guided quality optimization. The code for the metric can be found at https://github.com/gfxdisp/ColorVideoVDP.
VDP, color, image quality, video quality, color difference, perceptual metric
image quality, video quality, visual difference predictor, contrast sensitivity, visual metric
copyright: rightsretainedjournal: TOGjournalyear: 2024journalvolume: 43journalnumber: 4article: 129publicationmonth: 7doi: 10.1145/3658144conference: SIGGRAPH 2024 Technical Papers; July 28–Aug 1, 2024; Denver, USAbooktitle: SIGGRAPH 2024 Technical Papers (SIGGRAPH ’24 Technical Papers), Jul 28– Aug 1, 2024, Denver, USAccs: Computing methodologies Perceptionsubmissionid: papers_236
Figure 1. ColorVideoVDP predicts the visibility of distortions for a pair of test and reference videos (or images) as seen on a display with a provided specification. The predictions are represented as a single quality value in Just-Objectionable-Difference (JOD) units, a distortion map video, and a distogram, which visualizes the distortions over time, separately for each channel and spatial frequency band. See Fig. 25 for the content attribution.
[…]
1. Introduction
@@
Accurate reproduction of contrast and color is of central importance to the quality of displayed content. Achromatic artifacts stemming from graphics pipelines, such as visible blur or contrast loss, can be modelled by existing luminance-only metrics such as SSIM (Wang et al., 2003) or FovVideoVDP (Mantiuk et al., 2021), but color ones, such as chroma subsampling, cannot. On the other hand, color difference formulas, such as the popular CIEDE2000 (Sharma et al., 2005; CIE, 2018), do not model spatial or temporal aspects of vision, and as a consequence may ignore important aspects of an artifact, such as the spatial distribution or changes of color distortion over time. Notably, spatiotemporal color artifacts are especially problematic in modern display applications, in particular for emerging display technologies such as wide-color-gamut, virtual and augmented reality (XR) displays. The latter require novel architectural solutions, which lead to chromatic artifacts like color fringing caused by lens aberrations, or chromatic nonuniformity due to optical waveguides (see Fig. 1).
@@
This work presents ColorVideoVDP, a full-reference quality metric that models spatiotemporal achromatic and chromatic vision. The metric is built on novel psychophysical models of chromatic spatiotemporal contrast sensitivity and cross-channel contrast masking. Thanks to its psychophysical foundations, the metric accounts for the physical specification of a display (size, resolution, color characteristic) and viewing conditions (viewing distance and ambient light). This is the first video and image quality metric that explicitly models human spatiotemporal and chromatic vision simultaneously and is capable of modeling XR display artifacts. Our work builds on FovVideoVDP (Mantiuk et al., 2021) borrowing its general processing pipeline. The main novel contribution is integrating the spatio-chromatic model of near- and supra-threshold contrast perception, which lets us quantify the visibility of both achromatic and chromatic distortions. As chromatic distortions tend to reside in lower spatial frequencies, ColorVideoVDP also improves the modeling of low-frequency differences. Compared to FovVideoVDP, ColorVideoVDP is fully differentiable [^1] FovVideoVDP does not model the differences in low-frequency bands and, therefore, it cannot propagate low-frequency differences when used as a differentiable loss term. and much improves the accuracy of quality predictions on non-foveated image and video datasets.
@@
The key challenge of developing any new quality metric is its effective calibration and robust validation. To that end, and to ensure that the metric can provide reliable predictions for display applications, we collected a new XR-Display-Artifact-Video quality dataset (XR-DAVID) with 8 common display artifacts (Sec. 4). To ensure the diversity of distortion and content types, we combined our new XR-DAVID dataset with a large HDR/SDR image dataset UPIQ (Mikhailiuk et al., 2022). As the combined datasets consist of terabytes of data, calibration required a non-trivial mixture of end-to-end, and feature-space training. This effort allowed us to match and exceed the state-of-the-art results on unseen datasets (in a cross-dataset validation) and on the testing portion of the training datasets (cross-content validation). In Sec. 6, we demonstrate new applications of ColorVideoVDP, including analysis of chroma subsampling, display color tolerance specification, and quantifying observer metamerism variations on a target dataset.
@@
ColorVideoVDP lacks higher-level models of saliency or annoyance, resulting in lower accuracy when the semantic content has a strong influence on the quality judgments. It was not trained to predict accurate spatial distortion maps (Ye et al., 2019) (as no such data is available for video). ColorVideoVDP does not model the effect of glare (inter-ocular light scatter, found in HDR-VDP (Mantiuk et al., 2023)), gaze-contingent vision (found in FovVideoVDP (Mantiuk et al., 2021)), eye motion (Laird et al., 2006; Denes et al., 2020), or binocular vision (Didyk et al., 2011).
2. Related Work
Table 1. Quality/difference metrics and their capabilities. Columns indicate whether the metric models spatial vision, temporal vision, color vision, and whether it accounts for display geometry and photometry.
Metric Spatial Temporal Color Display model Approach
PSNR No No No No Signal quality
CIEDE2000 (CIE, 2018) No No Yes Yes Color difference formula
$\Delta\text{E}_{\text{ITP}}$ (ITU-R BT. 2124, 2019) No No Yes Yes Color difference formula
sCIELAB (Zhang and Wandell, 1997) Yes No Yes Yes CSF + Color difference formula
$\Delta\text{E}_{\text{ITP}}^{\text{S}}$ (Choudhury et al., 2021) Yes No Yes Yes CSF + Color difference formula
MS-SSIM (Wang et al., 2003) Yes No No No Multi-scale structural similarity
FSIMc (Zhang et al., 2011) Yes No Yes No Similarity of phase congruency and gradients
VSI (Zhang et al., 2014) Yes No Yes No Saliency + SSIM
LPIPS (Zhang et al., 2018) Yes No Yes No Difference of CNN features
FLIP (Andersson et al., 2020) Yes No Yes No CSF + Color difference + edge detectors
IQT (Cheon et al., 2021) Yes No Yes No CNN features + transformer autoencoder
STRRED (Soundararajan and Bovik, 2012) Yes Yes No No Entropy differences in wavelet subbands
VMAF (Li et al., 2016a) Yes Yes No No Features + SVR
FUNQUE (Venkataramanan et al., 2022) Yes Yes No No Wavelet decomposition + features + SVR
HDR-VDP-3 (Mantiuk et al., 2023) Yes No No Yes Psychophysical model
FovVideoVDP (Mantiuk et al., 2021) Yes Yes No Yes Psychophysical model
ColorVideoVDP (ours) Yes Yes Yes Yes Psychophysical model
@@ 2. Related Work
Perceived color differences for uniform patches can be predicted using one of the standard display formulas, such as CIE $\Delta\text{E}^{*}_{ab}$ or CIEDE2000 (Sharma et al., 2005; CIE, 2018). The standard CIE formulas, however, were not meant to predict differences for luminance below 1 cd/m2 or above the illuminance of 1 000 lux (CIE, 1993) (this translates to approximately 318 cd/m2 for a Lambertian white surface). $\Delta\text{E}_{\text{ITP}}$ (ITU-R BT. 2124, 2019) was proposed as a color difference formula for luminance levels found in wide-color-gamut high-dynamic-range television, and potentially intended for video content. It is unclear how to use the color difference formula with complex images and the color difference is typically computed per pixel and then averaged. Such treatment obviously ignores all spatial and temporal aspects of vision, which are partially addressed by the next group of metrics.
@@
The spatial component of vision was included in a spatial extension of the CIELAB difference formula — sCIELAB (Zhang and Wandell, 1997). The authors proposed to compute the CIE $\Delta\text{E}^{*}_{ab}$ color differences on images prefiltered by a contrast sensitivity function (CSF). Flip and HDR-Flip metrics (Andersson et al., 2020, 2021) improve on sCIELAB by employing a color difference formula that better quantifies large color differences. Both metrics also emphasize differences at edges, which tend to be more salient. Choudhury et al. proposed to compute color differences in the ITP color space (2021), which is more suitable for HDR color values. The strength of such spatial extensions is their simplicity. The main weakness is that such an application of the CSF is overly simplistic — it does not account for the changes in contrast sensitivity with luminance and does not account for supra-threshold vision (e.g. contrast masking and contrast constancy).
@@
The most popular image quality metrics, such as SSIM or MS-SSIM (Wang et al., 2003), do not attempt to explicitly model human vision, but, instead, they combine hand-crafted statistical measures that are likely to correlate with quality judgments. Although the early metrics, such as SSIM and MS-SSIM, operate only on the luma channel of the image, some later metrics, including FSIMc (Zhang et al., 2011) and VSI (Zhang et al., 2014), separate images into luma and two chroma channels, akin to the color space transforms used in video compression. Those metrics, however, do not account for the geometry of a display (e.g. resolution, size, viewing distance), nor for its photometry (e.g. peak brightness, black level). The latter shortcoming can be addressed by employing the Perceptually Uniform (PU) transform (Mantiuk and Azimi, 2021), which also extends these metrics to operate on high-dynamic-range images.
Zhang et al. (2018) observed that the activation layers of many convolution neural networks (CNNs) provide features that are well correlated with human judgments of image similarity. Their proposed metric, LPIPS, became very popular in computer vision, despite its underwhelming performance on image quality datasets (Ding et al., 2021). Prashnani et al. (2018) proposed training a CNN on triplets of patches, two distorted and one reference, with supervision based on the Bradley-Terry pairwise comparison model. Akin to LPIPS training, their PieAPP metric was trained on a large dataset of patches with pairwise comparison labels. Cheon et al. (2021) combined CNN features with transformer-based embeddings to regress quality scores. Their IQT metric was ranked first among 13 participants in the NTIRE 2021 perceptual image quality assessment challenge.
@@
Video quality metrics combine both spatial and temporal features. STRRED (Soundararajan and Bovik, 2012) compares the per-frame entropy of fitted local distributions of wavelet coefficients. This entropy is then weighted by local spatial and temporal variance. VMAF (Li et al., 2016a) combines two spatial features (VIF and DLM) with a mean absolute difference of consecutive frames, and then maps those into quality scores using a support vector regression (SVR). The success of VMAF motivated a series of fusion-based metrics, such as FUNQUE (Venkataramanan et al., 2022) and 3C-FUNQUE+ (Venkataramanan et al., 2023), which use wavelet transform modulated by the contrast sensitivity function as input to several individual predictors, which similarly as VMAF, are combined using SVR. However, all those metrics use simplified models of contrast sensitivity, which ignore the effect of luminance and do not model spatiotemporal vision. For example, the temporal processing of all of these state-of-the-art metrics considers just two consecutive frames, and it does not model the temporal characteristics of human vision. Our metric explicitly models achromatic and chromatic visual channels to address this shortcoming.
@@
The visual difference predictors (VDPs), such as DCTune (Watson, 1993), VDP (Daly, 1993), HDR-VDP-2 (Mantiuk et al., 2011) and HDR-VDP-3 (Mantiuk et al., 2023), explicitly model aspects of low-level human vision, such as contrast sensitivity and masking. The advantage of this approach is that these metrics are built on sound psychophysical models and generalize more easily to unseen conditions, such as displays of varying size, resolution, or peak luminance. While it is tempting to assume that this type of general modeling would perform worse or be more computationally expensive than metrics relying on hand-crafted features, as we will demonstrate in Sec. 5, this is not the case. Modern VDPs perform on par or better than metrics with hand-crafted features. When they are optimized to run on a GPU, they are as fast as feature-based metrics.
@@
Our ColorVideoVDP metric takes inspiration from and is based on similar building blocks as FovVideoVDP (Mantiuk et al., 2021) but with several important improvements. First, ColorVideoVDP models the visibility of chromatic (color) differences by employing a novel spatiotemporal-chromatic contrast sensitivity function (castleCSF) (Ashraf et al., 2024). This involves modeling both spatial and temporal chromatic vision. Second, we account for supra-threshold color differences so that the perceived magnitude of achromatic and chromatic contrast is properly mapped to the metric response. Third, while FovVideoVDP is insensitive to low-frequency distortions (below 1 cpd222cpd — cycles per visual degree — the measure of spatial frequency.), ColorVideoVDP correctly accounts for those. This lets us account for low-frequency display artifacts (e.g., waveguide non-uniformity) and use ColorVideoVDP as a differentiable loss function, which can propagate both low and high-frequency differences. Fourth, ColorVideoVDP models within-channel and cross-channel masking, where each of the four modeled channels (two achromatic and two chromatic) can be masked by the combination of contrast in the other channels. Finally, ColorVideoVDP is trained and tested on multiple image and video datasets, both SDR and HDR, including a novel dataset of display artifacts described in Sec. 4. Unlike FovVideoVDP, we do not model foveated vision due to the increased overhead of modeling interactions between eccentricity and color. While modeling foveated vision is necessary in some specific usage scenarios, such as foveated rendering, it is less useful in most general-use cases, such as video streaming, display engineering, and non-foveated algorithm design.
3. Color video visual difference predictor
@@
Figure 2. Processing stages of ColorVideoVDP. Test and reference images are first processed by the same pipeline: the display model maps pixel values to linear color (CIE 1931 XYZ color space), linear color is transformed to the opponent color space (DKL), the achromatic channel is decomposed into sustained (continuous lines) and transient (dashed lines) temporal channels, then each of those is decomposed into multiple spatial bands (Laplacian pyramid). The decomposed video/image goes into the contrast sensitivity and masking models, explained in more detail in Figure 6. The result of the masking model is pooled across all spatial bands, temporal and color channels, and finally regressed to a JOD score.